# Rectifying Conformity Scores for Better Conditional Coverage

### v1

@@ 1 Introduction
The widespread adoption of modern probabilistic and generative AI models across various applications highlights the need for reliable uncertainty quantification  (Gruber et al., 2023). In fact, while these models are highly flexible and can capture complex statistical dependencies, they can also generate predictions that are either unreliable or overly confident  (Nalisnick et al., 2018). Conformal prediction (CP; Vovk et al., 2005; Shafer & Vovk, 2008) is a robust framework for constructing distribution-free predictions with finite-sample validity guarantees; see (Angelopoulos et al., 2023, 2024). While most existing research has focused on univariate prediction tasks (Romano et al., 2019; Sesia & Romano, 2021; Rossellini et al., 2024), multivariate settings remain relatively underexplored. Conformal methods for multivariate regression include approaches that aggregate univariate prediction regions (Messoudi et al., 2021; Zhou et al., 2024), leverage generative models (Wang et al., 2023; Plassier et al., 2024a), or estimate conditional probability density functions (Izbicki et al., 2022).
Furthermore, classical CP methods ensure marginal validity but do not guarantee conditional validity, a stricter and more desirable property that tailors prediction regions to individual covariates. Unfortunately, prior research has demonstrated that constructing non-trivial prediction regions with exact conditional validity from finite samples is impossible without making additional assumptions about the data distribution (Vovk, 2012; Lei & Wasserman, 2014; Foygel Barber et al., 2021). As a result, efforts are often directed towards developing conformal methods that produce distribution-free prediction regions with marginal validity while aiming for the best possible approximate conditional validity (Gibbs et al., 2024).
A typical relaxation of exact conditional coverage in earlier work involves group-conditional guarantees (Ding et al., 2024; Jung et al., 2023), which provide coverage guarantees for a predefined set of groups. Another branch of work partitions the covariate space $\mathcal{X}$ into multiple regions and applies classical conformal prediction within each region (LeRoy & Zhao, 2021; Alaa et al., 2023; Kiyani et al., 2024). However, such partitioning based on the calibration set often leads to overly large prediction regions (Bian & Barber, 2023; Plassier et al., 2024b).
@@
Finally, several methods focus on the transformation of conformity scores (Deutschmann et al., 2023; Han et al., 2022; Dey et al., 2022; Izbicki et al., 2022; Dheur et al., 2024). These techniques adjust conformity scores to better approximate the conditional coverage. However, they usually require estimating the conditional distribution of conformity scores, which is both computationally intensive and difficult to perform accurately.
In this work, we introduce a new conformal prediction framework, Rectified Conformal Prediction (RCP), which extends the score transformation approach by applying transformations to arbitrary conformity scores. RCP is designed to improve conditional coverage while maintaining exact marginal coverage guarantees. By defining a new conformity score whose quantile (for a given coverage level) is independent of the covariates, RCP ensures that the conformal prediction procedure yields prediction sets that are both marginally and conditionally valid.
One of the key advantages of the RCP framework is its ability to produce multivariate prediction regions without requiring full modeling of the conditional distribution of conformity scores. Instead, RCP focuses on estimating specific quantiles of the score distribution that are critical for achieving the desired guarantees. This targeted estimation not only enhances computational efficiency but also makes the method versatile and applicable to a broad range of problems.
The main contributions of this work can be summarized as follows.
  • We introduce Rectified Conformal Prediction (RCP), a new conformal method designed to enhance conditional validity by refining conformity scores. Our method avoids the need to estimate the full conditional distribution of a multivariate response, relying instead on estimating only the conditional quantile of a univariate conformity score.
  • We provide a theoretical lower bound on the conditional coverage of the prediction sets generated by RCP. This conditional coverage is explicitly governed by the approximation error in estimating the conditional quantile of the conformity score distribution.
  • We evaluate our method on several benchmark datasets and compare it against state-of-the-art alternatives. Our results demonstrate improved performance, particularly in terms of conditional coverage metrics such as worst slab coverage (Romano et al., 2020) and conditional coverage error (Dheur et al., 2024).
@@ 2 Background
Consider a regression problem that aims to estimate a $d$-dimensional response vector $y\in\mathcal{Y}=\mathbb{R}^{d}$ based on a feature vector $x\in\mathcal{X}\subseteq\mathbb{R}^{p}$ to predict. We emphasize that one of the main goals of our work is the case of $d>1$. We denote by $\textup{P}_{X,Y}$ over $\mathcal{X}\times\mathcal{Y}$ the joint distribution of $(X,Y)$.
Construction of prediction regions for regression problems is often based on distributional regression that focuses on fully characterizing the conditional distribution of a response variable given a covariate (Klein, 2024). This approach improves uncertainty quantification and decision-making (Berger & Smith, 2019). From the conditional predictive distribution, prediction regions can be derived to capture values likely to occur with a given probability. However, these regions rely heavily on the predictive model’s quality, and poorly estimated models can result in unreliable predictions. In the following, we present split-conformal prediction (SCP; Papadopoulos et al., 2002), a computationally efficient variant of conformal prediction framework that allows to generate reliable prediction regions, even when the predictive model is misspecified or inaccurate.
@@
To do so, SCP relies on the so-called conformity score function, $V\colon\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$, which should assign larger value to worse agreement between $g(X)$ and $Y$. Let $\{(X_{k},Y_{k})\}_{k=1}^{n}$ be a calibration set, with $\mathcal{X}\subseteq\mathbb{R}^{p}$ and $\mathcal{Y}\subseteq\mathbb{R}^{d}$. SCP generates a prediction set $\mathcal{C}_{\alpha}(x)$ by computing an empirical quantile of the conformity scores $V(X_{k},Y_{k}),k=1,\dots,n$:
@@
where $\delta_{v}$ is the Dirac mass at $v$, and $Q_{1-\alpha}\left({\textup{P}}\right)$ denote the $(1-\alpha)$-quantile for any distribution P on $\mathbb{R}$.
@@
trivially satisfies conditional coverage (3) by the definition of conditional quantile $Q_{1-\alpha}(\textup{P}_{\mathbf{V}|X=x})$. However, exact conditional validity is not achievable within conformal prediction framework (Foygel Barber et al., 2021). In what follows we will present a new conformal prediction method that will achieve approximate conditional validity while satisfying exact marginal guarantees.
@@ 3 Rectified Conformal Prediction
A motivation for RCP.  Our RCP method is specifically designed to achieve both exact conformal marginal validity and improved approximate conditional coverage. To achieve this, RCP first constructs specially transformed (rectified) scores to enhance conditional coverage. To construct the rectified scores, it builds on the key observation that marginal and conditional coverage coincide precisely when the conditional $(1-\alpha)$-quantile of the conformity score is independent of the covariates. RCP then applies the SCP procedure to these rectified scores, ensuring the classical exact conformal marginal validity.
@@
Under 1, we denote by $\tilde{f}_{\varphi}^{-1}$ the inverse of the function $\tilde{f}_{\varphi}$, i.e. $\tilde{f}_{\varphi}^{-1}\circ\tilde{f}_{\varphi}(t)=t$, for all $t\in\mathbb{T}$. Let $\varphi\in\mathbb{R}$ be such that $\tilde{f}_{\varphi}$ is invertible (see 1). Set
@@
i.e. the $(1-\alpha)$ conditional quantile of the transformed score $\mathbf{V}_{\varphi}$ given $X=x$. We retrieve Example 1 with $f_{t}(v)=vt$, $\varphi=1$. In this case $\tilde{f}_{1}^{-1}(t)=t$ and $V_{1}(x,y)=V(x,y)$. Similarly, for Example 2, $f_{t}(v)=v+t$, $\varphi=0$. In such a case, $\tilde{f}_{0}^{-1}(t)=t$ and $V_{0}(x,y)=V(x,y)$.
@@
=$\quad\stackrel{{\scriptstyle(a)}}{{=}}\mathbb{P}(\mathbf{V}\leq\tilde{f}_{\varphi}(\tau_{\star}(X))\mid X=x)$
=$\quad\stackrel{{\scriptstyle(b)}}{{=}}\mathbb{P}(\mathbf{V}_{\varphi}\leq\tau_{\star}(X)\mid X=x)\stackrel{{\scriptstyle(c)}}{{\geq}}1-\alpha,$
@@ 4 Implementation of RCP
The RCP algorithm.   Rectified conformal prediction approach, as discussed above, requires a basic conformity score function $V$, a transformation function $f_{t}$, and a calibration dataset of $(n+m)$ points. A critical step in the RCP algorithm is estimating the conditional quantile $\widehat{\tau}(x)\approx Q_{1-\alpha}\bigl{(}\textup{P}_{\mathbf{V}_{\varphi}\mid X=x}\bigr{)}$, which we discuss in detail below. $\widehat{\tau}$ is learned on a separate part of calibration dataset composed of $m$ data points $\{(X_{k}^{\prime},Y_{k}^{\prime})\colon k=1,\dots,m\}$. Subsequently, RCP uses SCP with the rectified scores $\tilde{V}(x,y):=f_{\widehat{\tau}(x)}^{-1}\bigl{(}V(x,y)\bigr{)}$ instead of the basic scores $V(x,y)$. SCP is applied to the rectified scores computed on the second part of the calibration dataset: $\tilde{\mathbf{V}}_{k}=\tilde{V}(X_{k},Y_{k}),k=1,\dots,n$.
@@
[…]
@@
$\widehat{\tau}(x)=\arg\min_{\tau}\ \mathbb{E}\left[{\rho_{1-\alpha}\big{(}{V_{\varphi}(X,Y)-\tau(X)}\big{)}}\right],$ (12)
where the minimum is taken over the function $\tau\colon\mathcal{X}\to\mathbb{R}$ and $\rho_{1-\alpha}$ is the pinball loss  (Koenker & Bassett Jr, 1978; Koenker & Hallock, 2001): $\rho_{1-\alpha}(\tau)=(1-\alpha)\tau\mathds{1}_{\tau>0}-\alpha\tau\mathds{1}_{\tau\leq 0}$. In practice, the empirical quantile function is obtained $\widehat{\tau}$ by minimizing the empirical pinball loss:
$\widehat{\tau}\in\arg\min_{\tau\in\mathcal{C}}\frac{1}{m}\sum\nolimits_{k=1}^{m}\rho_{\alpha}\bigl{(}V_{\varphi}(X_{k}^{\prime},Y_{k}^{\prime})-\tau(X_{k}^{\prime})\bigr{)}+\lambda g(\tau),$ (13)
@@
Non-parametric methods have also been explored, mainly using kernel-based approaches; see (Christmann & Steinwart, 2008). In these cases, $g$ is a smoothness penalty (e.g a norm of an appropriately defined RKHS). More recently, Shen et al. (2024) introduced a penalized non-parametric approach to estimating the quantile regression process (QRP) using deep neural networks with rectifier quadratic unit (ReQU) activations. They provide non-asymptotic excess risk bounds for the estimated QRP under mild smoothness and regularity conditions.
@@
where $\{w_{k}(x)\}_{k=1}^{m}$ are positive weights; see (Bhattacharya & Gangopadhyay, 1990). For instance, we can set $w_{k}(x)=m^{-1}K_{h_{X}}(\|x-X_{k}^{\prime}\|)$, where for $h>0$, $K_{h}(\cdot)=h^{-1}K_{1}(h^{-1}\cdot)$ is a kernel function satisfying $\int K_{1}(x)\mathrm{d}x=1$, $\int xK_{1}(x)\mathrm{d}x=0$ and $\int x^{2}K_{1}(x)\mathrm{d}x<\infty$; $h_{X}$, the kernel bandwidth is tuned to balance bias and variance. With appropriate adaptive choice of $h(x)$, this approach can be shown to be asymptotically minimax over Hölder balls; see (Bhattacharya & Gangopadhyay, 1990; Spokoiny et al., 2013; Reiß et al., 2009).
5 Related works

### v2

@@ 1 Introduction
The widespread deployment of AI models emphasizes the need for reliable uncertainty quantification (Gruber et al., 2023). Although highly flexible in capturing complex statistical dependencies, these models can produce unreliable or overly confident predictions (Nalisnick et al., 2018). Conformal prediction (CP; Vovk et al. (2005); Shafer & Vovk (2008)) offers a robust, distribution-free framework for predictions with finite-sample validity guarantees (Angelopoulos et al., 2023, 2024).
Classical CP approaches guarantee marginal validity but fail to ensure the more desirable property of conditional validity, which customizes prediction regions to specific covariates. Prior studies have shown constructing meaningful prediction regions with exact conditional validity is infeasible without additional distributional assumptions (Vovk, 2012; Lei & Wasserman, 2014; Foygel Barber et al., 2021). Consequently, current research emphasizes developing conformal methods that maintain marginal validity and achieve approximate conditional validity (Colombo, 2024; Gibbs et al., 2025).
A typical relaxation of exact conditional coverage in earlier work involves group-conditional guarantees (Jung et al., 2023; Ding et al., 2024), which provide coverage guarantees for a predefined set of groups. Another branch of work partitions the covariate space $\mathcal{X}$ into multiple regions and applies CP within each set in the partition (LeRoy & Zhao, 2021; Alaa et al., 2023; Kiyani et al., 2024). However, such partitioning based on the calibration set often leads to overly large prediction regions (Bian & Barber, 2023; Plassier et al., 2024).
@@
Finally, several methods focus on the transformation of conformity scores (Han et al., 2022; Dey et al., 2022; Izbicki et al., 2022; Deutschmann et al., 2023; Dheur et al., 2024; Colombo, 2024). These techniques adjust conformity scores to better approximate the conditional coverage. However, they usually require estimating the conditional distribution of conformity scores, which is both computationally intensive and difficult to perform accurately.
In this paper, we propose a novel CP method, Rectified Conformal Prediction (RCP), extending normalized nonconformity scores; see, e.g., (Papadopoulos et al., 2008; Papadopoulos & Haralambous, 2011). RCP aims to enhance conditional coverage while preserving exact marginal coverage guarantees. By constructing a new conformity score whose quantile at a given coverage level is independent of covariates, RCP achieves both marginal and improved conditional validity.
A significant benefit of RCP is its capacity to generate prediction sets without fully modeling the conditional distribution of conformity scores. Instead, RCP concentrates on quantile regression to ensure approximate conditional coverage. The main contributions of this work can be summarized as follows.
  • We introduce Rectified Conformal Prediction (RCP), a new conformal method designed to enhance conditional validity by refining conformity scores (see Sections 3 and 4). The proposed method avoids the need to estimate the full conditional distribution of a multivariate response, relying instead on estimating only the conditional quantile of a univariate conformity score.
  • We provide a theoretical lower bound on the conditional coverage of the prediction sets generated by RCP (see Section 6). This conditional coverage is explicitly governed by the approximation error in estimating the conditional quantile of the conformity score distribution.
  • We evaluate our method on several benchmark datasets and compare it against state-of-the-art alternatives [^1] The code to reproduce main experiments is available at https://github.com/stat-ml/rcp (see Section 7). Our results demonstrate improved performance, particularly in terms of conditional coverage metrics such as worst slab coverage (Romano et al., 2020) and conditional coverage error (Dheur et al., 2024).
@@ 2 Background
Consider a regression problem that aims to estimate a $d$-dimensional response vector $y\in\mathcal{Y}=\mathbb{R}^{d}$ based on a feature vector $x\in\mathcal{X}\subseteq\mathbb{R}^{p}$ to predict. We denote by $\textup{P}_{X,Y}$ the joint distribution of $(X,Y)$ over $\mathcal{X}\times\mathcal{Y}$.
Construction of prediction regions for regression problems is often based on distributional regression that focuses on fully characterizing the conditional distribution of a response variable given a covariate (Klein, 2024). This approach improves uncertainty quantification and decision-making (Berger & Smith, 2019). From the conditional predictive distribution, prediction regions can be derived to capture values likely to occur with a given probability. However, these regions rely heavily on the predictive model’s quality, and poorly estimated models can result in unreliable predictions. In the following, we present split-conformal prediction (SCP; Papadopoulos et al., 2002), a computationally efficient variant of the conformal prediction framework that allows generating reliable prediction regions, even when the predictive model is misspecified or inaccurate.
@@
To do so, SCP relies on a conformity score function, $V\colon\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$, assigning larger value to worse agreement between $g(X)$ and $Y$. Let $\{(X_{k},Y_{k})\}_{k=1}^{n}$ be a calibration set, with $\mathcal{X}\subseteq\mathbb{R}^{p}$ and $\mathcal{Y}\subseteq\mathbb{R}^{d}$. SCP generates a prediction set $\mathcal{C}_{\alpha}(x)$ by computing an empirical quantile of the conformity scores $V(X_{k},Y_{k}),k=1,\dots,n$:
@@
where $\delta_{v}$ is the Dirac mass at $v$, and $Q_{1-\alpha}\left({\textup{P}}\right)$ denotes the $(1-\alpha)$-quantile for any distribution P on $\mathbb{R}$.
@@
trivially satisfies conditional coverage (2) by the definition of conditional quantile $Q_{1-\alpha}(\textup{P}_{\mathbf{V}|X=x})$. However, exact conditional validity is not achievable within conformal prediction framework (Vovk, 2012; Lei & Wasserman, 2014; Foygel Barber et al., 2021). In what follows we will present a new conformal prediction method that will achieve approximate conditional validity while satisfying exact marginal guarantees.
@@ 3 Rectified Conformal Prediction
A motivation for RCP.  Our RCP method is specifically designed to achieve both exact conformal marginal validity and approximate conditional coverage. To achieve this, RCP first constructs specially transformed (rectified) scores to enhance conditional coverage. To construct the rectified scores, it builds on the key observation that marginal and conditional coverage coincide precisely when the conditional $(1-\alpha)$-quantile of the conformity score is independent of the covariates. RCP then applies the SCP procedure to these rectified scores, ensuring the classical exact conformal marginal validity.
@@
Under 1, we denote by $\tilde{f}_{\varphi}^{-1}$ the inverse of the function $\tilde{f}_{\varphi}$, i.e., $\tilde{f}_{\varphi}^{-1}\circ\tilde{f}_{\varphi}(t)=t$, for all $t\in\mathbb{T}$. Let $\varphi\in\mathbb{R}$ be such that $\tilde{f}_{\varphi}$ is invertible (see 1). Set
@@
i.e., the $(1-\alpha)$ conditional quantile of the transformed score $\mathbf{V}_{\varphi}$ given $X=x$. We retrieve Example 1 with $f_{t}(v)=vt$, $\varphi=1$. In this case $\tilde{f}_{1}^{-1}(t)=t$ and $V_{1}(x,y)=V(x,y)$. Similarly, for Example 2, $f_{t}(v)=v+t$, $\varphi=0$. In such a case, $\tilde{f}_{0}^{-1}(t)=t$ and $V_{0}(x,y)=V(x,y)$.
@@
$\quad\stackrel{{\scriptstyle(a)}}{{=}}\mathbb{P}(\mathbf{V}\leq\tilde{f}_{\varphi}(\tau_{\star}(X))\mid X=x)$
$\quad\stackrel{{\scriptstyle(b)}}{{=}}\mathbb{P}(\mathbf{V}_{\varphi}\leq\tau_{\star}(X)\mid X=x)\stackrel{{\scriptstyle(c)}}{{\geq}}1-\alpha,$
@@ 4 Implementation of RCP
The RCP algorithm.  Rectified conformal prediction approach, as discussed above, requires a basic conformity score function $V$, a transformation function $f_{t}$, and a calibration dataset of $N=n+m$ points. A critical step in the RCP algorithm is estimating the conditional quantile $\widehat{\tau}(x)\approx Q_{1-\alpha}\bigl{(}\textup{P}_{\mathbf{V}_{\varphi}\mid X=x}\bigr{)}$, which we discuss in detail below. $\widehat{\tau}$ is learned on a separate part of calibration dataset composed of $m$ data points $\{(X_{k}^{\prime},Y_{k}^{\prime})\colon k=1,\dots,m\}$. Subsequently, RCP uses SCP with the rectified scores $\tilde{V}(x,y):=f_{\widehat{\tau}(x)}^{-1}\bigl{(}V(x,y)\bigr{)}$ instead of the basic scores $V(x,y)$. SCP is applied to the rectified scores computed on the second part of the calibration dataset: $\tilde{\mathbf{V}}_{k}=\tilde{V}(X_{k},Y_{k}),k=1,\dots,n$.
@@
Input: Calibration dataset $\mathcal{D}$, miscoverage level $\alpha$, conformity score function $V$, transformation function $f_{t}$, test input $x$.
$\triangleright$ Calibration Stage
 Split $\mathcal{D}$ into $\{(X_{k},Y_{k})\}_{k=1}^{n}$ and $\{(X_{k}^{\prime},Y_{k}^{\prime})\}_{k=1}^{m}$.
$\mathcal{D}_{\tau}\leftarrow\{(X_{k}^{\prime},V(X_{k}^{\prime},Y_{k}^{\prime}))\}_{k=1}^{m}$
$\widehat{q}_{1-\alpha}\leftarrow$ conditional quantile estimate on $\mathcal{D}_{\tau}$.
$\widehat{\tau}\leftarrow\tilde{f}_{\varphi}^{-1}\bigl{(}\widehat{q}_{1-\alpha}\bigr{)}$
for $k=1$ to $n$ do
  $\tilde{\mathbf{V}}_{k}\leftarrow f_{\widehat{\tau}(X^{\prime}_{k})}^{-1}\bigl{(}V(X_{k},Y_{k})\bigr{)}$.
end for
$k_{\alpha}\leftarrow\lceil(1-\alpha)(n+1)\rceil$.
$\tilde{\mathbf{V}}_{(k_{\alpha})}\leftarrow k_{\alpha}$-th smallest value in $\{\tilde{\mathbf{V}}_{k}\}_{k\in[n]}\cup\{+\infty\}$.
$\triangleright$ Test Stage
  $\mathcal{C}_{\alpha}(x)\leftarrow\bigr{\{}y\in\mathcal{Y}\colon f_{\widehat{\tau}(x)}^{-1}\bigl{(}V(x,y)\bigr{)}\leq\tilde{\mathbf{V}}_{(k_{\alpha})}\bigl{\}}$.
Output: $\mathcal{C}_{\alpha}(x)$.
@@
$\tau_{\star}(x)=\arg\min_{\tau}\ \mathbb{E}\left[{\rho_{1-\alpha}\big{(}{V_{\varphi}(X,Y)-\tau(X)}\big{)}}\right],$ (11)
where the minimum is taken over the function $\tau\colon\mathcal{X}\to\mathbb{R}$ and $\rho_{1-\alpha}$ is the pinball loss  (Koenker & Bassett Jr, 1978; Koenker & Hallock, 2001): $\rho_{1-\alpha}(\tau)=(1-\alpha)\tau\mathds{1}_{\tau>0}-\alpha\tau\mathds{1}_{\tau\leq 0}$. In practice, the empirical quantile function $\widehat{\tau}$ is obtained by minimizing the empirical pinball loss:
$\widehat{\tau}\in\arg\min_{\tau\in\mathcal{C}}\frac{1}{m}\sum_{k=1}^{m}\rho_{\alpha}\bigl{(}V_{\varphi}(X_{k}^{\prime},Y_{k}^{\prime})-\tau(X_{k}^{\prime})\bigr{)}+\lambda g(\tau),$ (12)
@@
Non-parametric methods have also been extensively explored, see, e.g., (Chernozhukov & Hansen, 2005; Chernozhukov et al., 2022). Takeuchi et al. (2006) introduced the kernel quantile regression (KQR) framework, formulating quantile regression as minimizing the pinball loss in an RKHS with Tikhonov (squared-norm) regularization. It established some of the first theoretical guarantees for RKHS-based quantile models, deriving finite-sample generalization error bounds using Rademacher complexity; these results were later improved in (Li et al., 2007)
@@
where $\{w_{k}(x)\}_{k=1}^{m}$ are positive weights; see (Bhattacharya & Gangopadhyay, 1990). For instance, we can set $w_{k}(x)=m^{-1}K_{h_{X}}(\|x-X_{k}^{\prime}\|)$, where for $h>0$, $K_{h}(\cdot)=h^{-1}K_{1}(h^{-1}\cdot)$ is a kernel function satisfying $\int K_{1}(x)\mathrm{d}x=1$, $\int xK_{1}(x)\mathrm{d}x=0$ and $\int x^{2}K_{1}(x)\mathrm{d}x<\infty$; $h_{X}$, the kernel bandwidth is tuned to balance bias and variance. With appropriate adaptive choice of $h(x)$, this approach can be shown to be asymptotically minimax over Hölder balls; see (Bhattacharya & Gangopadhyay, 1990; Spokoiny et al., 2013; Reiß et al., 2009). More recently, Shen et al. (2024) introduced a penalized non-parametric approach to estimating the quantile regression process (QRP) using deep neural networks with rectifier quadratic unit (ReQU) activations. Shen et al. (2024) derives upper bounds on the mean-squared error for quantile regression using deep ReQU networks, depending only on the approximation error and network. The bounds are shown to be tight for broad function classes (e.g., Hölder compositions, Besov spaces), implying that ReQU neural networks achieve minimax-optimal convergence rates for conditional quantile estimation. Notably, the theory requires minimal assumptions and holds even for heavy-tailed error distributions.
5 Related Work